# Incorporating Geo-Diverse Knowledge into Prompting for Increased Geographical Robustness in Object Recognition

### v1

@@ 3.1 Geographical Knowledge Probing
For CountryLLM, we use the geo-specific set:
[…]
The argmax of $s$ with respect to $c$ is taken as the prediction. Due to averaging over descriptor scores, not every descriptor needs to activate for an inference to be correct. The model therefore can account for diverse features of objects within a geography. These descriptors effectively serve as complements to CLIP’s default knowledge of class names.
Combining knowledge. Our third method, CountryInPrompt+LLM, combines both CLIP’s internal knowledge and LLM external knowledge. The prompt template $\boldsymbol{t}_{c,d}^{\text{CountryInPrompt+LLM}}$ is “a photo of a/an $<$object$>$ in $<$country$>$ which (is/has/etc.) $<$descriptor$>$”.
3.2 Regularizing Soft Prompts via Geo Knowledge
Adaptation scenario. In practice, one may want to adapt VLMs for a downstream task. To update model parameters, one promising strategy is soft textual prompting. It is parameter-efficient [52] and avoids feature distortion unlike finetuning [24]. However, it involves learning context parameters that directly change the class text embeddings used in inference. Learning context on a dataset with limited diversity (e.g. just Europe) may thus tailor these class representations to the region and overfit. To investigate cross-geography generalization when using soft prompting, we pose a domain generalization scenario where we aim to learn from a high-resource source set of countries and generalize to a target set of unseen countries at inference time. A method that performs well in this setting could provide a viable alternative to few-shot target training when acquiring target data for training is not feasible.
[…]
${\boldsymbol{t}_{c}=[\text{V}]_{1}[\text{V}]_{2}...[\text{V}]_{M}[\text{CLASS${}_{c}$}]}$ (3)
The respective class text embedding $\boldsymbol{w}_{c}$ is produced as $h(\boldsymbol{t}_{c})$ after forwarding the prompt through the text encoder. Learning proceeds by minimizing cross-entropy, for image $k$ with features $\boldsymbol{f}_{k}$, using ground-truth source labels $y_{k,c}$ and temperature $\tau$:
@@ 3.2 Regularizing Soft Prompts via Geo Knowledge
Geography knowledge regularization (gkr). We minimize the cosine distance of normalized class embedding $\boldsymbol{w_{c}}$ and overall target class knowledge $\boldsymbol{k}_{c}^{tgt}$:
@@
Geo knowledge ensemble. To define $\boldsymbol{k}_{c}^{tgt}$, we identify that a model may be deployed in various geographies. Therefore, we define a target geography set $\mathcal{G}_{t}$ (e.g. in Africa, Asia, Americas), which can be practically thought of as countries a model may be deployed in that are not in the training set $\mathcal{D}$ (e.g. Europe). Then for each geography $g$ in $\mathcal{G}_{t}$, we define the corresponding class knowledge $\boldsymbol{k}_{c}^{g}$ as:
@@
This is defined analogously for CountryInPrompt and CountryLLM. The final regularization target $\boldsymbol{k}_{c}^{tgt}$ for class $c$ aggregates the set’s geography-specific knowledge:
@@
Table 1: Zero-shot CLIP with descriptive knowledge prompts, top-1/3 balanced accuracy (Acc) on DollarStreet. Our strategies to capture CLIP’s internal country knowledge (CountryInPrompt), external LLM country knowledge (CountryLLM), and their combination (CountryInPrompt+LLM), improve the zero-shot CLIP baseline (prompt “a photo of a”), especially on Africa (exemplified in light blue) and Asia; gains in green, drops in red. Our strategies also outperform the GeneralLLM [30] baseline.
@@
Table 2: Regularizing soft prompts with country knowledge, top-1 bal. acc. on DollarStreet. We emphasize our regularization aims to improve target performance on unseen geographies, rather than source (gray, italicized). Gains/drops vs. the best of soft prompt baselines (shaded). Our CountryInPrompt+LLM achieves notable gains in target, especially on Africa (light blue). Methods use 16 shots per class.
@@ 4 Experimental Setup
Datasets. DollarStreet [11] has 38,479 images of household objects across regions (Africa, Americas - South/Central/North, Asia, Europe) and various incomes. The classes may represent abstract concepts (e.g. most loved toys), so we narrow focus to 95 object classes (23,114 images). We merge especially close categories (light sources by bed/in living room) and ignore multi-label examples. For zero-shot inference, the entire set is used. For training, source is Europe, and target is Americas, Asia, Africa. The source uses a train-val-test split of 64%-16%-20% (stratified based on class); evaluation is on all data from target continents. To set up $\boldsymbol{k}_{c}^{tgt}$, the 49 target countries make up $\mathcal{G}_{t}$. GeoNet [21] contains three benchmarks including GeoImNet, comprised of 171,692 images across 600 object categories from the USA (source) and 78,358 images from the same number of categories in Asia (target). We use existing train-test splits for soft prompt training. To set up $\boldsymbol{k}_{c}^{tgt}$, given the relatively large number of categories and inference costs of davinci-003, $\mathcal{G}_{t}$ is set to contain the top 10 most frequently represented countries in Asia.
Baselines. We compare our geography knowledge regularization method to CoOp [52], CoCoOp [51] and KgCoOp [47]. For zero-shot inference, we compare to CLIP (default prompts) and the classification via description of [30].
@@ 5.1 Zero-shot CLIP Inference with Geo Knowledge
We gauge the zero-shot effectiveness of three prompting strategies: (1) CountryInPrompt (including country names in prompts to probe CLIP’s internal knowledge), (2) CountryLLM (gathering external descriptive knowledge of objects with davinci-003), and (3) CountryInPrompt+LLM (using country names and LLM knowledge). We compare to the method of [30] (GeneralLLM) and default zero-shot CLIP with manual prompts (i.e., “a photo of a $<$object$>$”). Results on DollarStreet are shown in Table 1.
Including country names in object prompts improves geographical robustness, especially in Africa and Asia. This observation is supported by gains for CountryInPrompt vs. Zero-Shot CLIP, especially in Africa and Asia (up to +5.4 and +2.6 top-1 gains for RN50, resp.) CLIP is thus shown to possess useful knowledge of country-specific object categories that can be effectively probed with a simple prompt. In Americas/Europe, adding country names leads to gains with RN50, but slight drops with ViT-B/16 and ViT-B/32. We reason that CLIP’s default prompts may be already well-aligned to countries in these regions for those encoders. We posit that adding country names is most effective when there is a large shift between CLIP’s default prompts and a country’s specific object representation.
Prompting with country-specific descriptive knowledge from LLMs outperforms general object knowledge [30], as we see from CountryLLM’s larger gains than GeneralLLM over default CLIP in almost all encoders, regions, and metrics. The largest top-1 difference is shown with ViT-B/32 (in Total, 52.6% for CountryLLM vs. 51.4% for GeneralLLM). Using top-3 accuracy, we observe the following differences for CountryLLM/GeneralLLM in Total (74.6/73.0 for ViT-B/32, 78.8/77.9 for ViT-B/16, 70.0/68.6 for RN50). These suggest that default non-country-specific knowledge may be biased and less adequate for various countries. The gains of CountryLLM vs. Zero-Shot CLIP are generally largest on Africa and Asia, as countries in these regions may have greater shifts vs. the default prompts, but CountryLLM also performs well on Europe.
There are complementary benefits to using CLIP’s internal and external LLM geo knowledge. This observation is supported by CountryInPrompt+LLM, the combination of CountryInPrompt and CountryLLM, achieving the best Total top-1 performance for every encoder. The Total gains vs. default CLIP are as large as +3.4 (RN50). While CLIP has internal knowledge of country-specific categories, it may be incomplete due to limited representation in the image-text training corpus. Adding LLM world knowledge, trained on a different, purely textual corpus, may address the gaps. CountryInPrompt+LLM is notably the top setting in 3 of 4 regions for each encoder in terms of top-1 accuracy.
@@ 5.2 Soft Prompting
We next evaluate geographical knowledge regularization (Sec. 3.2), our method to improve target performance by ensuring that soft prompts do not overfit class text representations to a source dataset with limited geographic representativeness (e.g. only data from Europe). We compare regularization with ensembles of CountryInPrompt, CountryLLM, and CountryInPrompt+LLM prompts vs. state-of-the-art soft prompting methods in Table 2/3.
Regularizing soft prompts with target geographical knowledge reduces overfitting to source geographies. Our method effectively improves the ability of CLIP, with prompts trained only on images from Europe, to generalize to unseen countries. This observation is supported by the gains in Total Target accuracy for CountryInPrompt Reg, CountryLLM Reg, and CountryInPrompt+LLM Reg on DollarStreet (+1.1, +0.8, and +1.6 over the best of the soft prompt baselines for ViT-B/16, respectively). The improvements are especially notable in Africa: our CountryInPrompt+LLM Reg method achieves +2.8 for ViT-B/16 and +1.8 for RN50 vs. all soft prompt baselines. The effectiveness of our strategy extends to GeoNet in Table 3: accuracy on target increases by +1.3 for ViT-B/16 and +1.4 for RN50 over the best soft prompt baseline. Our combined strategy works best on the target. These results show the value in incorporating both internal CLIP knowledge and external LLM knowledge to address concept shift in the optimization process. Since regularization prevents overfitting and potentially optimal performance on the source, we naturally observe source drops for CountryInPrompt and CountryLLM in Tables 2 and 3. However, CountryInPrompt+LLM in Table 2 even offers gains on source.
@@
CIPReg 50.6 -1.4 47.6 +0.7
LLMReg 51.8 -0.2 47.4 +0.5
CIP+LLMReg 51.1 -0.9 48.3 +1.4
Table 3: Regularizing soft prompts with geography knowledge, top-1 bal. accuracy on GeoNet. The regularization method accomplishes our goal to increase target performance in GeoNet’s USA-to-Asia transfer setting. CIP = CountryInPrompt, LLM = CountryLLM, CIP+LLM= CountryInPrompt+LLM.
@@
Geographic regularization helps significantly on difficult classes. As certain objects may be especially sensitive to geographical domain shift, we break down classwise performance on DollarStreet in Table 4, using a stratification of class difficulty based on default soft prompting performance (CoOp). Our CountryInPrompt+LLM strategy achieves significant gains on the classes most difficult with respect to the CoOp baseline. In particular, gains of +8.7% in balanced accuracy are achieved for classes with $<$40% baseline accuracy, while the highest achieved by KgCoOp is 4.1%. The 13 classes in this subset are: bowls/plates, snacks, clothes, dish racks, dish soap, hand/body soap, detergent, makeup, mosquito protections, piercings, refrigerators, showers, trash. The DollarStreet classes with greatest improvement, independent of original CoOp accuracy, are: piercings, clothes, homes, medication, and refrigerators (all at least 14% improved over CoOp). In GeoNet, dome, goby (fish), eland (antelope), and gloriosa (flower) have at least 20 samples and greater than 30% improvement. A total of 64/95 classes in DollarStreet and 209/344 in GeoNet improve with our method vs. CoOp, showing broad coverage.
@@
Performance by income. DollarStreet provides estimated monthly income of the household in which an image was captured. We use the delineation of low, medium, and high-income buckets from [13]. Compared to CoOp/KgCoOp, our CountryInPrompt+LLM achieves gains of +2.5/+3.4 in low, +2.4/+1.5 in medium, and +2.1/+0.7 in high. Thus our method especially improves performance in low-income areas, and works well for geo shifts even at a finer granularity of shifts (household income). We include the table in supp.
@@ 5.3 Further Analysis
Are LLM descriptors correlated with key country statistics? For a pair of countries, we measure the distance between each class embedding, and take the average distance as one value. The other value is the absolute difference between statistics for those countries (from [2, 42], e.g. difference in avg. yearly temperature). We compute the correlation between these two values over every unique country pair in DollarStreet, showing results in Table 5. We find that the strongest correlation across each prompt type is with HDI, which summarizes human development. It is notable that factors like yearly temperature and precipitation also show moderate correlations, indicating a potential role of climate. Future work may further explore how object differences present with respect to these factors.
@@
Figure 5: UMAP [29] plot for CountryLLM and the category homes in DollarStreet. Country-specific descriptors are often close to those of other countries intra-continent, likely due to similar weather, environment and/or economic conditions.
Descriptor topics. We show a UMAP [29] visualization comparing CountryLLM text embeddings for the category homes across geographies in Fig. 5. Countries tend to group by continent, showing the class representations may capture similarity in features like climate and/or economics. We examine a few topics mentioned in the CountryLLM descriptors for homes. While “stone” is represented across continents, “mud” and “bright colors” are mentioned predominantly in Africa, “balcony” in Europe and Asia, and “flower” in Europe and America. We show more in supp.
Success and failure examples. We provide examples of CountryLLM vs. GeneralLLM in Fig. 4 using DollarStreet. The model captures geographical descriptive knowledge like “sandy colored stucco walls” for homes in Jordan, a feature which may be less common for Western homes. Sometimes the model may be overly attentive to attributes, leading to confusion (e.g. choosing rug over bed).
@@ 6 Conclusion
We show that the combination of geography-specific descriptive knowledge from LLMs and CLIP’s learned geographical knowledge yield cross-geography recognition improvements when effectively incorporated into zero-shot inference and regularization for soft prompts learned on easy-to-access source data. Gains are most significant on data from Africa and for the most difficult classes.
Limitations and ethical considerations. Our method is limited to the knowledge in CLIP and LLMs, which can be incorrect (as LLMs are known to hallucinate), imprecise, or biased to the worldview the Internet provides.
Future work. While we show utility in leveraging LLM and VLM knowledge, we advocate for future work that ensures factuality and cultural accuracy within this knowledge, including to prevent exaggerated differences between countries. We will explore additional knowledge to expand utility to various regions and groups, and communicate with diverse communities to gather knowledge and data to ensure fairness in AI technology use. We will also consider how to individually address the various components of geographical domain shift (e.g. due to material, style, and context).

### v2

@@ 3.2 Regularizing Soft Prompts via Geo Knowledge
The respective class text embedding $\boldsymbol{w}_{c}$ is produced as $h(\boldsymbol{t}_{c})$, forwarding the prompt through the text encoder. Learning proceeds by minimization of cross-entropy, for image $k$ with features $\boldsymbol{f}_{k}$, using ground-truth source labels $y_{k,c}$ and temperature $\tau$:
@@
Geography knowledge regularization (gkr). We minimize the cosine distance of normalized class embedding $\boldsymbol{w_{c}}$ and overall target class knowledge $\boldsymbol{k}_{c}^{tgt}$, over all $c$:
@@
Geo knowledge ensemble. To define $\boldsymbol{k}_{c}^{tgt}$, we identify that a model may be deployed in various locations. Therefore, we define a target geography set $\mathcal{G}_{t}$, which can practically be thought of as the countries that a model may be deployed in that are not in the training set $\mathcal{D}$ (e.g. Africa, Asia, Americas in $\mathcal{G}_{t}$ if only Europe in $\mathcal{D}$). Then for each geography $g$ in $\mathcal{G}_{t}$, we define the corresponding class knowledge $\boldsymbol{k}_{c}^{g}$ as:
@@
This is defined analogously for CountryInPrompt and CountryLLM. The final regularization target $\boldsymbol{k}_{c}^{tgt}$ for class $c$ aggregates the set’s geographical knowledge:
@@
Table 1: Zero-shot CLIP inference with descriptive knowledge prompts, top-1/3 balanced accuracy (Acc) on DollarStreet. Strategies to capture CLIP’s internal country knowledge (CountryInPrompt), external LLM country knowledge (CountryLLM), and their combination (CountryInPrompt+LLM), often improve vs. the zero-shot CLIP baseline (prompt “a photo of a/an $<$object$>$”), especially on Africa and Asia; gains in green, drops in red. CountryLLM notably outperforms the GeneralLLM [30] baseline.
@@
Table 2: Regularizing soft prompts with geographical knowledge, top-1 bal. acc. on DollarStreet. We emphasize that our regularization aims to improve target performance, rather than source (gray, italicized). Gains/drops are shown vs. the best of soft prompt baselines (shaded). CountryInPrompt+LLM Reg achieves notable gains in target, especially on Africa. Methods use 16 shots per class.
@@ 4 Experimental Setup
Datasets. We use DollarStreet [11], which has 38,479 images of household objects across regions (Africa, South/Central/North America, Asia, Europe) and incomes. The classes may represent abstract concepts (e.g. most loved toys), so we narrow focus to 95 object classes. We merge especially close categories (light sources by bed/in living room) and ignore multi-label examples, resulting in 23,114 total images. For zero-shot inference, the entire set is used. For training, the source is Europe, and the target is Americas, Asia, and Africa. 20% of source data, stratified based on class proportions, is heldout for testing; target evaluation is on all data from target continents. To set up $\boldsymbol{k}_{c}^{tgt}$, the 49 target countries in DollarStreet make up $\mathcal{G}_{t}$. We also use the GeoImNet benchmark of GeoNet [21], comprised of 171,692 images across 600 objects from the USA (source) and 78,358 images across the same number in Asia (target). We use existing train-test splits for soft prompt training. For GeoNet, given the relatively large number of categories and inference costs of davinci-003, $\boldsymbol{k}_{c}^{tgt}$ and $\mathcal{G}_{t}$ use the top 10 most frequent countries in the GeoNet set.
Baselines. We evaluate geography knowledge regularization vs. CoOp [52], CoCoOp [51] and KgCoOp [47]. For zero-shot inference, we evaluate CLIP with default prompts and the classification via description method of [30].
@@ 5.1 Zero-shot CLIP Inference with Geo Knowledge
We gauge the effectiveness of three zero-shot strategies: (1) CountryInPrompt (including countries in prompts to probe CLIP’s knowledge), (2) CountryLLM (gathering descriptive knowledge of objects with davinci-003), and (3) CountryInPrompt+LLM (using country names and LLM knowledge). We compare to [30] (GeneralLLM) and CLIP with manual prompts (i.e. “a photo of a/an $<$object$>$”). Results on DollarStreet are shown in Table 1.
Including country names in prompts can improve object recognition, especially in Africa and Asia. This observation is supported by gains for CountryInPrompt vs. Zero-Shot CLIP, especially in Africa and Asia (up to +5.4 and +2.6 top-1 gains for RN50, resp.). Such differences may occur as country-specific context can align representations closer to these regions, while default prompts do not adequately capture objects around the world (esp. from non-Western regions). In Americas/Europe, adding country names leads to gains with RN50, but slight drops with ViT-B/16 and ViT-B/32. We reason that CLIP’s default prompts may be already well-aligned to countries in these regions for those encoders due to overrepresentation in training.
Prompting with country-specific descriptive knowledge from LLMs outperforms general object knowledge. We observe this from CountryLLM’s larger gains over default CLIP than GeneralLLM’s for almost all encoders, regions, and metrics. The largest top-1 difference is with ViT-B/32 (in Total, 52.6% for CountryLLM vs. 51.4% for GeneralLLM). In top-3 accuracy, the differences for CountryLLM/GeneralLLM in Total are 74.6/73.0 for ViT-B/32, 78.8/77.9 for ViT-B/16, 70.0/68.6 for RN50. These suggest that default non-country-specific knowledge is less adequate for various countries. The gains of CountryLLM vs. Zero-Shot CLIP are generally largest on Africa and Asia, as countries in these regions may have greater shifts vs. the default prompts, but CountryLLM also performs well on Europe. LLM description in general is less effective in Americas, though Americas has a large proportion of USA images, for which default CLIP may be well-aligned.
There are complementary effects when using CLIP’s internal and external LLM geo knowledge. This observation is supported by CountryInPrompt+LLM, the combination of CountryInPrompt and CountryLLM, achieving the best Total top-1 performance for every encoder. The Total gains vs. default CLIP are as large as +3.4 (RN50). While CLIP has internal knowledge of country-specific categories, it may be incomplete and imprecise due to limited representation in the image-text training corpus. Adding LLM knowledge, trained on a purely textual corpus, may address some gaps. CountryInPrompt+LLM is notably the top setting in 3/4 regions for each encoder in top-1 accuracy.
@@ 5.2 Soft Prompting
We next evaluate geography knowledge regularization (Sec. 3.2), our method to improve target performance by ensuring that soft prompts do not overfit class text representations to a source dataset with limited geographical representativeness (e.g. only data from Europe). We compare regularization with ensembles of CountryInPrompt, CountryLLM, and CountryInPrompt+LLM prompts vs. state-of-the-art soft prompting methods in Tables 2/3.
Regularizing soft prompts with target geographical knowledge reduces overfitting to source geographies. Our method effectively improves the ability of CLIP, with prompts trained only on images from Europe, to generalize to target countries. This observation is supported by Total Target gains for CountryInPrompt, CountryLLM, and CountryInPrompt+LLM Reg on DollarStreet (+1.1/0.8/1.6 over the best soft prompt baseline for ViT-B/16). Improvements are notable in Africa: CountryInPrompt+LLM achieves +2.8 for ViT-B/16 and +1.8 for RN50. The effectiveness extends to GeoNet in Table 3: target gains are +1.3 for ViT-B/16 and +1.4 for RN50. The combined strategy works best on target, showing the value of incorporating descriptive knowledge. Since regularization prevents overfitting and potentially optimal source performance, we naturally observe source drops for CountryInPrompt and CountryLLM in Tables 2/3. However, CountryInPrompt+LLM in Table 2 even offers source gains. It is also notable that there are small drops in Americas for RN50, but upon inspection, countries in North America overall have a -0.9 drop, while ones in Central/South America have a +0.8 gain. These results concur with our hypothesis that CLIP is already aligned to countries like the USA. More is in supp., along with experiments varying the source and ensemble.
@@
CIPReg 50.6 -1.2 47.6 +0.7
LLMReg 51.8 0.0 47.4 +0.5
CIP+LLMReg 51.1 -0.7 48.3 +1.4
Table 3: Regularizing soft prompts with geographical knowledge, top-1 bal. accuracy on GeoNet. The regularization method accomplishes our goal to increase target performance in GeoNet’s USA-to-Asia transfer setting. CIP = CountryInPrompt, LLM = CountryLLM, CIP+LLM = CountryInPrompt+LLM.
@@
Regularization helps significantly on difficult classes. As certain objects may be especially sensitive to geographical domain shift, we break down classwise performance on DollarStreet in Table 4, using a stratification of class difficulty based on default soft prompting performance (CoOp). The CountryInPrompt+LLM strategy achieves significant gains on the classes most difficult with respect to the CoOp baseline. In particular, gains of +8.7% in balanced accuracy are achieved for classes with $<$40% baseline recall, while the highest achieved by KgCoOp is 4.1%. Example classes in this subset are snacks, clothes, and makeup. The DollarStreet classes with greatest improvement, independent of original CoOp accuracy are: piercings, clothes, homes, medication, and refrigerators (all at least +14% over CoOp). In GeoNet, dome, goby (fish), eland (antelope), and gloriosa (flower) have $>$20 samples and $>$30% improvement. A total of 64/95 classes in DollarStreet and 209/344 in GeoNet improve vs. CoOp, showing broad coverage.
@@
Performance by income. DollarStreet provides estimated monthly income of the household in which an image was captured. We evaluate with the delineation of low, medium, and high-income buckets from [13]. Compared to CoOp/KgCoOp, CountryInPrompt+LLM gains are +2.5/+3.4 in low, +2.4/+1.5 in medium, and +2.1/+0.7 in high. Thus our method especially improves in low-income areas, though it helps across levels. The table is in supp.
@@ 5.3 Further Analysis
Are descriptions correlated with key country statistics? For a pair of countries, we compute two values. We measure the distance between each class embedding and take the average overall distance as one value. We also take the absolute difference between statistics for those countries (from [2, 42], e.g. difference in avg. yearly temperature) as the other value. We compute the correlation between these two values over every unique country pair in DollarStreet, showing results in Table 5. We find that the strongest correlation across each prompt type is with HDI, which summarizes human development. It is notable that factors like yearly temperature and precipitation also show moderate correlations, indicating a potential role of climate. Future work may further explore how object differences present with respect to these factors. It will also be critical to ensure that differences between countries are representative and not exaggerated in embeddings.
@@
Figure 5: UMAP [29] plot for CountryLLM and the category homes in DollarStreet. Country-specific descriptors are often close to those of other countries intra-continent, likely due to similar weather, environment, and/or economic conditions.
Descriptor topics. We show a UMAP [29] visualization comparing CountryLLM text embeddings for the category homes across geographies in Fig. 5. Countries tend to group by continent, showing the representations may capture similarity in features like climate and/or economics. We examine a few topics mentioned in the CountryLLM descriptors for homes. While “stone” is described across continents, “bright colors” and “mud” are mentioned mostly in Africa, and “balcony” in Europe and Asia. We show more in supp.
Success and failure examples. We provide examples of CountryLLM vs. GeneralLLM in Fig. 4 on DollarStreet. The model captures geographical descriptive knowledge like “sandy colored stucco walls” for homes in Jordan, a feature which may be less common for Western homes. Sometimes the model may be too attentive to attributes, leading to confusion (e.g. choosing rug over bed). Future work that enhances alignment in VLMs can likely improve results.
@@ 6 Conclusion
In this work, we bring attention to how various strategies to prompt CLIP affect recognition performance across geographies. In addition, through soft prompting with descriptive knowledge, we provide a mechanism to achieve a more geo-generalizable set of class representations across regions. Our work is only a first step in this important area.
Limitations and ethical considerations. While our method’s proof of concept is demonstrated in a positive effort to debias CLIP’s default representations through diversity, due to the biased worldview of the Internet, CLIP’s representations are likely inadequate, exaggerated, and/or not fully representative for some countries. While we expect quality LLM knowledge to guide better representations, LLM knowledge can also be incorrect (e.g., through hallucination), imprecise, or biased.
Future work. For the above reasons, our future efforts aim to ensure more representative VLM/LLM knowledge. We strongly advocate for the community to seek communication with diverse groups within all countries (i.e. to capture areas that range from low to high income) to ensure better representation and fairness in AI technology use. There are notable continent-level disparities to still improve upon.
Acknowledgement. This work was supported by National Science Foundation Grants No. 2006885 and 2329992.